# Efficient 3D Molecular Generation with Flow Matching and Scale Optimal Transport

### v1

@@ Abstract
Generative models for 3D drug design have gained prominence recently for their potential to design ligands directly within protein pockets. Current approaches, however, often suffer from very slow sampling times or generate molecules with poor chemical validity. Addressing these limitations, we propose Semla, a scalable E(3)-equivariant message passing architecture. We further introduce a molecular generation model, MolFlow, which is trained using flow matching along with scale optimal transport, a novel extension of equivariant optimal transport. Our model produces state-of-the-art results on benchmark datasets with just 100 sampling steps. Crucially, MolFlow samples high quality molecules with as few as 20 steps, corresponding to a two order-of-magnitude speed-up compared to state-of-the-art, without sacrificing performance. Furthermore, we highlight limitations of current evaluation methods for 3D generation and propose new benchmark metrics for unconditional molecular generators. Finally, using these new metrics, we compare our model’s ability to generate high quality samples against current approaches and further demonstrate MolFlow’s strong performance.
@@ 1 Introduction
In this work we tackle this problem from two directions. Firstly, we introduce a novel equivariant architecture for 3D molecular generation which exhibits significantly better efficiency and scalability than existing approaches. Even with 3 times as many parameters as the current state-of-the-art, our model processes a batch of molecules more than twice as quickly. Secondly, we introduce a novel type of optimal transport, scale optimal transport, which extends equivariant optimal transport [23, 47] to account for different sizes of molecules. Using our model architecture, we then train an unconditional molecular generator, MolFlow, using flow matching with scale optimal transport. With only 100 sampling steps our model outperforms all existing approaches on two benchmark datasets. We further show that with as few as 20 sampling steps MolFlow’s performance is comparable to the existing state-of-the-art while producing a 2-orders-of-magnitude speedup in sampling times. Finally, we highlight issues with current evaluation metrics and introduce two new benchmark metrics for unconditional molecular generation to combat these.
@@
  • MolFlow – A state-of-the-art molecular generator trained using flow matching, which provides up to a 70-fold improvement in sampling time compared to existing approaches.
@@
Figure 1: An overview of our approach. Firstly, prior coordinate noise is sampled according to the respective scale of the molecules (atom and bond type noise is not shown for brevity). Our trained model, MolFlow, transports this noise to the data distribution to generate a molecular sample.
@@ 4 Flow Matching for Molecular Generation
To assess the ability of Semla to model distributions with $E(3)\times S_{N}$ symmetry we apply conditional flow matching [28, 2, 1, 30] with optimal transport to create a generative model for molecules, which we refer to as MolFlow. In this section we outline the training and sampling procedure for MolFlow with molecular structures. In the following we represent each molecule by a tuple $z=(\mathbf{x},\mathbf{a},\mathbf{b},\mathbf{c})$ of coordinates, atom types, bond types and formal charges, respectively.
Training MolFlow
As shown in existing conditional flow matching frameworks [28, 2, 10], training proceeds by firstly sampling: noise $z_{0}\sim p_{\mathrm{noise}}(z_{0})$; data $z_{1}\sim p_{\mathrm{data}}(z_{1})$; and a time $t\in\left[0,1\right]$, and using these to sample from the time-dependent conditional flow $z_{t}\sim p_{t|1}(z|z_{0},z_{1})$. In practice, we apply scale optimal transport (OT) and sample noise from the scaled distribution $p_{\mathrm{noise}}^{\pi}(z_{0}|n)$, where $n$ is the number of atoms. We discuss this further, along with the OT transformation applied to the noise during training below. Previous work on molecular structure generation has found it beneficial to train models to predict data directly rather than noise [26] or a vector field [48]. MolFlow is therefore trained to learn a distribution $p_{1|t}^{\theta}(z_{1}|z_{t})$ which predicts clean data from interpolated data using a Semla model with parameters $\theta$. Different conditional flows are used for continuous and discrete data. For coordinates, we apply a linear interpolation with the addition of a small amount of Gaussian noise. Atom and bond types are sampled according to the recently proposed discrete flow models (DFM) framework [10]. The joint molecular interpolation is therefore given as follows:
@@
$\mathcal{L}_{\mathrm{MolFlow}}=\lambda_{x}\mathcal{L}_{\mathrm{MSE}}(\mathbf{\tilde{x}}_{1},\mathbf{x}_{1})+\lambda_{a}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{a}}_{1},\mathbf{a}_{1})+\lambda_{b}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{b}}_{1},\mathbf{b}_{1})+\lambda_{c}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{c}}_{1},\mathbf{c}_{1})$ (12)
We also make use of self-conditioning, which was originally proposed for diffusion models as way of reusing the model’s previous prediction when sampling [11]. To create a self-conditioned MolFlow model, we adopt the same training procedure as HarmonicFlow [48]. We provide further details on this, along with the hyperparameters used for MolFlow, in Appendix C.
@@
Once we have trained a MolFlow model, molecules can be generated by, firstly, sampling noise $z_{0}\sim p_{\mathrm{noise}}^{\pi}(z_{0}|n)$, and then integrating the ODE corresponding to the conditional flow $p_{t|1}$ from $t=0$ to $t=1$. For coordinates, the vector field corresponding to our choice of conditional flow is given by $\mathbf{\tilde{x}}_{1}-\mathbf{x}_{0}$ where $\mathbf{\tilde{x}}_{1}$ is sampled from $p_{1|t}^{\theta}$ as shown above. We then apply an Euler solver to integrate the ODE with step sizes $\Delta t$ as follows: $\mathbf{x}_{t+\Delta t}=\mathbf{x}_{t}+\Delta t(\mathbf{\tilde{x}}_{1}-\mathbf{x}_{0})$. We refer readers to DFM [10] for the sampling procedure for atom and bond types. In practice, we found that taking logarithmically spaced steps resulted in better performance than using constant step sizes.
@@ 5 Experiments
In this section we provide results on benchmark 3D molecular generation tasks and compare the performance of our model to existing state-of-the-art approaches. In Appendix B, we provide results on ablation experiments which show how different components of the training, such as scale OT and self conditioning, affect performance on molecular generation tasks. We also provide samples from a MolFlow model trained on GEOM Drugs in Appendix E.
@@
Two benchmark datasets, QM9 [37] and GEOM Drugs [4], are used to assess MolFlow’s abilities as an unconditional molecular generator. QM9 is a complete enumeration of molecules with up to 9 heavy atoms which satisfy a set of pre-defined constraints. On average, molecules in QM9 contain approximately 18 atoms. GEOM Drugs, on the other hand, is a dataset of larger, drug-like molecules with an average size of around 44 atoms. For both datasets we use the same data splits as MiDi and EQGAT-diff. To improve training times, however, we discard molecules with more than 72 atoms from the GEOM Drugs training set. This corresponds to about 1% of the training data. Validation and test sets are left unchanged. All metrics for MolFlow presented below are calculated by, firstly, sampling molecule sizes from the test set, and then generating molecules with the sampled number of atoms using the model, as outlined above.
@@
MolFlow (Ours) 99.9±${}_{\pm\text{0.0}}$ 99.6±${}_{\pm\text{0.06}}$ 99.4±${}_{\pm\text{0.12}}$ 97.5±${}_{\pm\text{0.20}}$ 100
@@
We compare MolFlow to a number of recently proposed models for 3D molecular generation, including both models which infer bonds based on coordinates and models which generate bonds directly. EDM [19], GCDM [33], MUDiff [20] and GFMDiff [58] are all diffusion-based models which infer bonds from atom positions. We also compare to EquiFM [47] which uses flow-matching along with equivariant optimal transport to generate atom types and coordinates; bonds are then inferred based on these. Finally, we also compare to MiDi [56] and EQGAT-diff [26] which learn molecular bonds and which we regard as the current state-of-the-art. We use standard benchmark evaluation metrics: atom stability; molecule stability; validity; uniqueness; and novelty, which have been thoroughly described in previous works. We also provide a full description of these metrics in Appendix D. Additionally, for each baseline model we provide the number of function evaluations (NFE) required to sample one batch of molecules. These were taken from the respective publications, if available, or using the default values in the official codebases.
@@
Table 1 compares the performance of MolFlow with existing approaches. Following [55, 19] we do not provide novelty scores on QM9. MolFlow is trained for 300 epochs on QM9 on a single Nvidia A100 GPU, and we provide further training details in Appendix C. From the table we can see that MolFlow matches or exceeds all models on all metrics other than uniqueness, despite using 5 times fewer sampling steps than MiDi and EQGAT-diff. Our model also significantly outperforms EquiFM, the only other flow-matching-based model in the table. We also note that since EquiFM uses an adaptive ODE solver the number of function evaluations in the table is an average and could be significantly higher for larger, more complex molecules.
Since QM9 contains only very small molecules GEOM Drugs serves a more useful benchmark for distinguishing model performance. Table 2 compares our model to MiDi and EQGAT-diff on this dataset. We do not include models which infer bonds since these models typically perform very poorly on larger molecules and often do not provide results for all evaluation metrics. Once again we see that MolFlow outperforms existing state-of-the-art models on all key evaluation metrics, while generating only marginally fewer novel molecules. In addition to requiring significantly fewer evaluation steps our model also requires much less compute for training. MolFlow trains for 200 epochs on a single Nvidia A100 GPU, compared to 800 epochs with 4 GPUs for EQGAT-diff.
@@
MolFlow (Ours) 99.8±${}_{\pm\text{0.01}}$ 97.7±${}_{\pm\text{0.07}}$ 95.2±${}_{\pm\text{0.22}}$ 100.0±${}_{\pm\text{0.0}}$ 99.6±${}_{\pm\text{0.09}}$ 100
@@
To further improve the efficiency of our model we tried varying the number of ODE integration steps and we provide a one-to-one comparison of these results with EQGAT-diff in Table 3. We also measure the time required by each model to sample 5000 molecules. For EQGAT-diff we measure the sample time using the given evaluation code with a batch size of 75, which was the largest multiple of 25 we could consistently fit on an Nvidia A100 40GB GPU. In the table we can see a stark difference in runtime between the two models. MolFlow provides comparable performance to EQGAT-diff with as few as 20 sampling steps, corresponding to a two order of magnitude speed-up in sampling time. Much of this improvement is due to the very small number of sampling steps, which we hypothesise is due to the use of optimal transport, since this aims to minimise the cost of transporting probability mass between distributions. However, we further see a two-fold improvement in the time per forward pass, which demonstrates the importance of designing efficient equivariant architectures for molecular generation.
@@
Table 3: Comparison between EQGAT-diff and MolFlow with different numbers of sampling steps. Molecule stability and validity are given as percentages, energy and strain energy are measured in $\mathrm{kcal}\cdot\mathrm{mol}^{-1}$, NFE refers to the number of function evaluations required to sample one batch of molecules and sample time is measured by the number of seconds required to generate 5000 molecules, averaged over 5 runs.
@@
MolFlow20 94.8±${}_{\pm\text{0.28}}$ 93.7±${}_{\pm\text{0.36}}$ 168.7±${}_{\pm\text{3.1}}$ ${}_{\pm\text{3.1}}$ 20 167
MolFlow50 97.2±${}_{\pm\text{0.25}}$ 95.2±${}_{\pm\text{0.40}}$ 141.1±${}_{\pm\text{2.1}}$ 102.0±${}_{\pm\text{2.1}}$ 50 412
MolFlow100 97.7±${}_{\pm\text{0.07}}$ 95.2±${}_{\pm\text{0.22}}$ ${}_{\pm\text{2.1}}$ 88.9±${}_{\pm\text{1.3}}$ 100 821
@@
Table 3 shows that MolFlow produces molecules with better energies and strain energies than EQGAT-diff, despite the significantly faster sampling time. Notably, however, molecules generated by EQGAT-diff have lower minimised energies than MolFlow, suggesting that their model is better at finding molecular conformations which have lower energy minima, while our model is better at producing lower strain energies. In ablation experiments in Appendix B we also show that the use of scale optimal transport is a crucial component for generating low energy conformations with our model.
@@ 7 Conclusion
In this work we have presented Semla, a novel equivariant message passing architecture which exhibits much better efficiency and scalability than existing approaches for unconditional molecular generation. Additionally, we introduced a novel type of optimal transport (OT), scale OT, which extends equivariant OT but noticeably reduces the transport cost. We then trained a Semla model for 3D molecular generation using flow matching with scale OT. Our model, MolFlow, achieves state-of-the-art results on molecular generation benchmarks and, crucially, is able to generate high quality molecules with only 20 sampling steps. We also highlighted issues with current molecular evaluation metrics and proposed the use of energy and strain for evaluating the quality of generated molecular conformations.
While we believe our model has made significant progress in solving key challenges for 3D molecular generators, many challenges remain. Firstly, the energies of the molecules generated by MolFlow are still significantly higher than that of the dataset; generating molecular coordinates with very high fidelity remains a problem for these models. Including further inductive biases or fine-tuning against an energy model could be an avenue to improve this in future work [34, 43, 57]. Additionally, while MolFlow has shown significant efficiency improvements over existing methods, it still uses a full-connected message passing component, limiting it’s scalability to larger molecular systems. We leave the further enhancement of the scalability of Semla to future work. We believe our model makes crucial step towards the practical application of 3D molecular generators, although we leave the integration of MolFlow into drug discovery workflows, either through RL-based fine-tuning or protein pocket conditioned generation, to future work. In its current form our model has the potential to be used for malicious activities, however, we believe the potential positive contributions of this model, helping to design new therapeutics, for example, strongly outweigh the negatives.

### v2

@@ Abstract
Generative models for 3D drug design have gained prominence recently for their potential to design ligands directly within protein pockets. Current approaches, however, often suffer from very slow sampling times or generate molecules with poor chemical validity. Addressing these limitations, we propose Semla, a scalable E(3)-equivariant message passing architecture. We further introduce a molecular generation model, SemlaFlow, which is trained using flow matching along with scale optimal transport, a novel extension of equivariant optimal transport. Our model produces state-of-the-art results on benchmark datasets with just 100 sampling steps. Crucially, SemlaFlow samples high quality molecules with as few as 20 steps, corresponding to a two order-of-magnitude speed-up compared to state-of-the-art, without sacrificing performance. Furthermore, we highlight limitations of current evaluation methods for 3D generation and propose new benchmark metrics for unconditional molecular generators. Finally, using these new metrics, we compare our model’s ability to generate high quality samples against current approaches and further demonstrate SemlaFlow’s strong performance.
@@ 1 Introduction
In this work we tackle this problem from two directions. Firstly, we introduce a novel equivariant architecture for 3D molecular generation which exhibits significantly better efficiency and scalability than existing approaches. Even with 3 times as many parameters as the current state-of-the-art, our model processes a batch of molecules more than twice as quickly. Secondly, we introduce a novel type of optimal transport, scale optimal transport, which extends equivariant optimal transport [23, 47] to account for different sizes of molecules. Using our model architecture, we then train an unconditional molecular generator, SemlaFlow, using flow matching with scale optimal transport. With only 100 sampling steps our model outperforms all existing approaches on two benchmark datasets. We further show that with as few as 20 sampling steps SemlaFlow’s performance is comparable to the existing state-of-the-art while producing a 2-orders-of-magnitude speedup in sampling times. Finally, we highlight issues with current evaluation metrics and introduce two new benchmark metrics for unconditional molecular generation to combat these.
@@
  • SemlaFlow – A state-of-the-art molecular generator trained using flow matching, which provides up to a 70-fold improvement in sampling time compared to existing approaches.
@@
Figure 1: An overview of our approach. Firstly, prior coordinate noise is sampled according to the respective scale of the molecules (atom and bond type noise is not shown for brevity). Our trained model, SemlaFlow, transports this noise to the data distribution to generate a molecular sample.
@@ 4 Flow Matching for Molecular Generation
To assess the ability of Semla to model distributions with $E(3)\times S_{N}$ symmetry we apply conditional flow matching [28, 2, 1, 30] with optimal transport to create a generative model for molecules, which we refer to as SemlaFlow. In this section we outline the training and sampling procedure for SemlaFlow with molecular structures. In the following we represent each molecule by a tuple $z=(\mathbf{x},\mathbf{a},\mathbf{b},\mathbf{c})$ of coordinates, atom types, bond types and formal charges, respectively.
Training SemlaFlow
As shown in existing conditional flow matching frameworks [28, 2, 10], training proceeds by firstly sampling: noise $z_{0}\sim p_{\mathrm{noise}}(z_{0})$; data $z_{1}\sim p_{\mathrm{data}}(z_{1})$; and a time $t\in\left[0,1\right]$, and using these to sample from the time-dependent conditional flow $z_{t}\sim p_{t|1}(z|z_{0},z_{1})$. In practice, we apply scale optimal transport (OT) and sample noise from the scaled distribution $p_{\mathrm{noise}}^{\pi}(z_{0}|n)$, where $n$ is the number of atoms. We discuss this further, along with the OT transformation applied to the noise during training below. Previous work on molecular structure generation has found it beneficial to train models to predict data directly rather than noise [26] or a vector field [48]. SemlaFlow is therefore trained to learn a distribution $p_{1|t}^{\theta}(z_{1}|z_{t})$ which predicts clean data from interpolated data using a Semla model with parameters $\theta$. Different conditional flows are used for continuous and discrete data. For coordinates, we apply a linear interpolation with the addition of a small amount of Gaussian noise. Atom and bond types are sampled according to the recently proposed discrete flow models (DFM) framework [10]. The joint molecular interpolation is therefore given as follows:
@@
$\mathcal{L}_{\mathrm{SemlaFlow}}=\lambda_{x}\mathcal{L}_{\mathrm{MSE}}(\mathbf{\tilde{x}}_{1},\mathbf{x}_{1})+\lambda_{a}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{a}}_{1},\mathbf{a}_{1})+\lambda_{b}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{b}}_{1},\mathbf{b}_{1})+\lambda_{c}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{c}}_{1},\mathbf{c}_{1})$ (12)
We also make use of self-conditioning, which was originally proposed for diffusion models as way of reusing the model’s previous prediction when sampling [11]. To create a self-conditioned SemlaFlow model, we adopt the same training procedure as HarmonicFlow [48]. We provide further details on this, along with the hyperparameters used for SemlaFlow, in Appendix C.
@@
Once we have trained a SemlaFlow model, molecules can be generated by, firstly, sampling noise $z_{0}\sim p_{\mathrm{noise}}^{\pi}(z_{0}|n)$, and then integrating the ODE corresponding to the conditional flow $p_{t|1}$ from $t=0$ to $t=1$. For coordinates, the vector field corresponding to our choice of conditional flow is given by $\mathbf{\tilde{x}}_{1}-\mathbf{x}_{0}$ where $\mathbf{\tilde{x}}_{1}$ is sampled from $p_{1|t}^{\theta}$ as shown above. We then apply an Euler solver to integrate the ODE with step sizes $\Delta t$ as follows: $\mathbf{x}_{t+\Delta t}=\mathbf{x}_{t}+\Delta t(\mathbf{\tilde{x}}_{1}-\mathbf{x}_{0})$. We refer readers to DFM [10] for the sampling procedure for atom and bond types. In practice, we found that taking logarithmically spaced steps resulted in better performance than using constant step sizes.
@@ 5 Experiments
In this section we provide results on benchmark 3D molecular generation tasks and compare the performance of our model to existing state-of-the-art approaches. In Appendix B, we provide results on ablation experiments which show how different components of the training, such as scale OT and self conditioning, affect performance on molecular generation tasks. We also provide samples from a SemlaFlow model trained on GEOM Drugs in Appendix E.
@@
Two benchmark datasets, QM9 [37] and GEOM Drugs [4], are used to assess SemlaFlow’s abilities as an unconditional molecular generator. QM9 is a complete enumeration of molecules with up to 9 heavy atoms which satisfy a set of pre-defined constraints. On average, molecules in QM9 contain approximately 18 atoms. GEOM Drugs, on the other hand, is a dataset of larger, drug-like molecules with an average size of around 44 atoms. For both datasets we use the same data splits as MiDi and EQGAT-diff. To improve training times, however, we discard molecules with more than 72 atoms from the GEOM Drugs training set. This corresponds to about 1% of the training data. Validation and test sets are left unchanged. All metrics for SemlaFlow presented below are calculated by, firstly, sampling molecule sizes from the test set, and then generating molecules with the sampled number of atoms using the model, as outlined above.
@@
SemlaFlow (Ours) 99.9±${}_{\pm\text{0.0}}$ 99.6±${}_{\pm\text{0.06}}$ 99.4±${}_{\pm\text{0.12}}$ 97.5±${}_{\pm\text{0.20}}$ 100
@@
We compare SemlaFlow to a number of recently proposed models for 3D molecular generation, including both models which infer bonds based on coordinates and models which generate bonds directly. EDM [19], GCDM [33], MUDiff [20] and GFMDiff [58] are all diffusion-based models which infer bonds from atom positions. We also compare to EquiFM [47] which uses flow-matching along with equivariant optimal transport to generate atom types and coordinates; bonds are then inferred based on these. Finally, we also compare to MiDi [56] and EQGAT-diff [26] which learn molecular bonds and which we regard as the current state-of-the-art. We use standard benchmark evaluation metrics: atom stability; molecule stability; validity; uniqueness; and novelty, which have been thoroughly described in previous works. We also provide a full description of these metrics in Appendix D. Additionally, for each baseline model we provide the number of function evaluations (NFE) required to sample one batch of molecules. These were taken from the respective publications, if available, or using the default values in the official codebases.
@@
Table 1 compares the performance of SemlaFlow with existing approaches. Following [55, 19] we do not provide novelty scores on QM9. SemlaFlow is trained for 300 epochs on QM9 on a single Nvidia A100 GPU, and we provide further training details in Appendix C. From the table we can see that SemlaFlow matches or exceeds all models on all metrics other than uniqueness, despite using 5 times fewer sampling steps than MiDi and EQGAT-diff. Our model also significantly outperforms EquiFM, the only other flow-matching-based model in the table. We also note that since EquiFM uses an adaptive ODE solver the number of function evaluations in the table is an average and could be significantly higher for larger, more complex molecules.
Since QM9 contains only very small molecules GEOM Drugs serves a more useful benchmark for distinguishing model performance. Table 2 compares our model to MiDi and EQGAT-diff on this dataset. We do not include models which infer bonds since these models typically perform very poorly on larger molecules and often do not provide results for all evaluation metrics. Once again we see that SemlaFlow outperforms existing state-of-the-art models on all key evaluation metrics, while generating only marginally fewer novel molecules. In addition to requiring significantly fewer evaluation steps our model also requires much less compute for training. SemlaFlow trains for 200 epochs on a single Nvidia A100 GPU, compared to 800 epochs with 4 GPUs for EQGAT-diff.
@@
SemlaFlow (Ours) 99.8±${}_{\pm\text{0.01}}$ 97.7±${}_{\pm\text{0.07}}$ 95.2±${}_{\pm\text{0.22}}$ 100.0±${}_{\pm\text{0.0}}$ 99.6±${}_{\pm\text{0.09}}$ 100
@@
To further improve the efficiency of our model we tried varying the number of ODE integration steps and we provide a one-to-one comparison of these results with EQGAT-diff in Table 3. We also measure the time required by each model to sample 5000 molecules. For EQGAT-diff we measure the sample time using the given evaluation code with a batch size of 75, which was the largest multiple of 25 we could consistently fit on an Nvidia A100 40GB GPU. In the table we can see a stark difference in runtime between the two models. SemlaFlow provides comparable performance to EQGAT-diff with as few as 20 sampling steps, corresponding to a two order of magnitude speed-up in sampling time. Much of this improvement is due to the very small number of sampling steps, which we hypothesise is due to the use of optimal transport, since this aims to minimise the cost of transporting probability mass between distributions. However, we further see a two-fold improvement in the time per forward pass, which demonstrates the importance of designing efficient equivariant architectures for molecular generation.
@@
Table 3: Comparison between EQGAT-diff and SemlaFlow with different numbers of sampling steps. Molecule stability and validity are given as percentages, energy and strain energy are measured in $\mathrm{kcal}\cdot\mathrm{mol}^{-1}$, NFE refers to the number of function evaluations required to sample one batch of molecules and sample time is measured by the number of seconds required to generate 5000 molecules, averaged over 5 runs.
@@
SemlaFlow20 94.8±${}_{\pm\text{0.28}}$ 93.7±${}_{\pm\text{0.36}}$ 168.7±${}_{\pm\text{3.1}}$ ${}_{\pm\text{3.1}}$ 20 167
SemlaFlow50 97.2±${}_{\pm\text{0.25}}$ 95.2±${}_{\pm\text{0.40}}$ 141.1±${}_{\pm\text{2.1}}$ 102.0±${}_{\pm\text{2.1}}$ 50 412
SemlaFlow100 97.7±${}_{\pm\text{0.07}}$ 95.2±${}_{\pm\text{0.22}}$ ${}_{\pm\text{2.1}}$ 88.9±${}_{\pm\text{1.3}}$ 100 821
@@
Table 3 shows that SemlaFlow produces molecules with better energies and strain energies than EQGAT-diff, despite the significantly faster sampling time. Notably, however, molecules generated by EQGAT-diff have lower minimised energies than SemlaFlow, suggesting that their model is better at finding molecular conformations which have lower energy minima, while our model is better at producing lower strain energies. In ablation experiments in Appendix B we also show that the use of scale optimal transport is a crucial component for generating low energy conformations with our model.
@@ 7 Conclusion
In this work we have presented Semla, a novel equivariant message passing architecture which exhibits much better efficiency and scalability than existing approaches for unconditional molecular generation. Additionally, we introduced a novel type of optimal transport (OT), scale OT, which extends equivariant OT but noticeably reduces the transport cost. We then trained a Semla model for 3D molecular generation using flow matching with scale OT. Our model, SemlaFlow, achieves state-of-the-art results on molecular generation benchmarks and, crucially, is able to generate high quality molecules with only 20 sampling steps. We also highlighted issues with current molecular evaluation metrics and proposed the use of energy and strain for evaluating the quality of generated molecular conformations.
While we believe our model has made significant progress in solving key challenges for 3D molecular generators, many challenges remain. Firstly, the energies of the molecules generated by SemlaFlow are still significantly higher than that of the dataset; generating molecular coordinates with very high fidelity remains a problem for these models. Including further inductive biases or fine-tuning against an energy model could be an avenue to improve this in future work [34, 43, 57]. Additionally, while SemlaFlow has shown significant efficiency improvements over existing methods, it still uses a full-connected message passing component, limiting it’s scalability to larger molecular systems. We leave the further enhancement of the scalability of Semla to future work. We believe our model makes crucial step towards the practical application of 3D molecular generators, although we leave the integration of SemlaFlow into drug discovery workflows, either through RL-based fine-tuning or protein pocket conditioned generation, to future work. In its current form our model has the potential to be used for malicious activities, however, we believe the potential positive contributions of this model, helping to design new therapeutics, for example, strongly outweigh the negatives.